OpenStudy (anonymous):

write an equation of the line containing the given point and perpendicular to the given line. (0,5); 4x+3y=7

OpenStudy (anonymous):

slope of line = -4/3 perpendicular line has slope of 3/4 use y - y1 = m(x-x1) so y - 5 = (3/4) x - 0) y = (3/4)x + 5 or 4y = 3x + 20

OpenStudy (anonymous):

you ok with that?

OpenStudy (anonymous):

perfect! thank you so much!

OpenStudy (anonymous):

yeah tats good

OpenStudy (anonymous):

just understand if a line is perpendicular its slope is -1/m ok

OpenStudy (anonymous):

wrt to the line which it is perpendicular to